The frequency of a recessive allele in a large randomly mating population is 0.2.
What is the frequency of:
(a) The dominant allele?
(b) The heterozygote
1
Expert's answer
2018-05-24T07:40:07-0400
a. q=0.2, and p+q=1. Thus, the frequency of dominant allele is 1-0.2=0.8
b. The frequency of heterozygote is 2*p*q. It is 2*0.8*0.2 = 0.32